One of the most powerful tools for improving success rates in introductory computing courses is the incorporation of context – a theme that pervades the computing lectures, assignments, and examples which relates the content to a concrete application domain. Contextualized computing education has even allowed us to be successful with challenging audiences, such as the non-technical major. In this talk, we review why Georgia Tech has chosen to teach serious computer science to every student on campus, and then discuss research findings from several schools on the benefits and costs of contextualized computing education.